OpenStudy (anonymous):

A sound wave has a wavelength of 0.450 meters. If its speed in cold air is 330 meters/second, what is the wave's frequency? 0.890 hertz 100 hertz 230 hertz 733 hertz

OpenStudy (michele_laino):

the relationship between frequency \(f\), wavelength \(\lambda\) and speed \(v\) is: \(\lambda \cdot f= v\)

OpenStudy (michele_laino):

so: \[f = \frac{v}{\lambda } = \frac{{330}}{{0.450}} = ...?\]